A circular hole of 15 cm diameter is cut on a copper plate at a temperature of 30 celsius . Find the area of hole at 150 degree celsius.
1
Expert's answer
2019-11-06T09:17:19-0500
"A'=A(1+\alpha(T_2-T_1))=0.25\pi d^2(1+\alpha(T_2-T_1))"

"A'=0.25\pi (15)^2(1+1.7\cdot10^{-5}(150-30))=177.075\ cm^2"